OpenStudy (anonymous):

solve for K, show your work 13k + 5 = k - 7

OpenStudy (lgbasallote):

first... subtract k from both sides...what do you get?

OpenStudy (anonymous):

13k - 1k = 7 - 5 12k = 2

OpenStudy (lgbasallote):

carefull... it's -7 - 5... not 7- 5

OpenStudy (lgbasallote):

careful*

OpenStudy (lgbasallote):

so you have 12k = -7 - 5 what does this simplify to?

OpenStudy (anonymous):

ohh, it inherits its sign okay.

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

12k = -12 k = -1

OpenStudy (anonymous):

13k - k = -7 - 5 What do you get?

OpenStudy (lgbasallote):

right